OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 ANSWERS BELOW PLS HELP

OpenStudy (anonymous):

slope = -5; y-intercept = (0, -6) slope = five thirds; y-intercept = (0, 2) slope = -three fifths; y-intercept = (0, 2) slope = five thirds; y-intercept = (0, -6)

OpenStudy (anonymous):

B. slope= -5/3; y-intercept= (0,2)